You are here: Foswiki>ARGOS Web>Commissioning>DiaryRun15>DayTime20170302 (03 Mar 2017, TommasoMazzoni)Edit Attach

Power up issues

When we power up the SX LGSW the Pupil motor RED Y axis was not homing. We unmount the LGSW panel and we push by hand the stage and at this point the motor complete the home procedure (http://aowiki.arcetri.astro.it/ARGOSPublic/NSA12NotHoming#Procedure_to_recover_from_a_motor_stuck_in_the_fully_elongated_position). When we power up again the system the Pupil motor YELLOW Y axis was not homing. We repeat the procedure and then it works.

When we power up the LGSW from the arbitrator we got an error because the WFS camera controller gone in the "ErrorState". Here the log from the WFS camera controller

2017-03-02 16:35:56.635241214 ^[[31mERROR^[[0m WfsCamera sx-lgsw:9781.7fb7d2db9700 nobody WfsCameraDriver.cpp:71 void argos::wfs_camera::WfsCameraDriver::logAndThrowBasdaWfsCameraException(const std::string&) [^[[1m[State] 'StartPnCCD()' ignored. Cannot turn cooling on. Check statuses.^[[0m]

The PSU comminicating led in the GUI was red (Debug Statuses tab). we restart the kontron and start the pnCCD from the gui and it works.

Pockel cell transmission

TNs Plot
SX: ['20170302_175423', ..., '20170302_175546'] 20170302_175423_pcTransmission.jpg
SX: ['20170302_224804', ..., '20170302_224923'] 20170302_224804_pcTransmission.jpg

In 5 hours the performance of the BLUE and the RED pockel cells decrease, in particular the right part of the BLUE goes from more the 100 (suppression factor) to 20.

SX AO calibration

SX HEXA 4.72 2.18 -0.3 66 68
With HBS on we see more vibration then without HBS.

HBS on, TN 20170302_190439: 1236nm of Tip/Tilt, all Open Loop

HBS off, TN 20170302_190814: 303nm of Tip/Tilt, all Open Loop

We work with HBS off.

150 modes in OL
MH LGSW IM LGSW REC COMBO LGSW REC
20170302_183150 20170302_192450 20170302_194100 20170302_215400

150 modes in CL on LGSW REC 20170302_194100
MH LGSW IM LGSW REC COMBO LGSW REC
20170302_183150 20170302_201317 20170302_221300 20170302_221300
The new modal History use the old modal amplitudes divide by a factor 2.

We test the new LGSW rec and we compare with an old one from October 2016. When we try to increase the gain the subaps on the top right start to oscillate. 0.8" disturbance.
TN AO Gain AO CL TT Res. WF [nm rms] HO Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20170302_220725 0.5,0.5,0.5 T 151.034 64.6863 20161013_153800 -1 F F
20170302_220740 0.5,0.5,0.5 T 149.052 63.693 20161013_153800 -1 F F
20170302_220751 0.5,0.5,0.5 T 143.001 62.8341 20161013_153800 -1 F F
20170302_221115 0.5,0.4,0.3 T 157.292 73.0444 20170302_215400 -1 F F
20170302_221130 0.5,0.4,0.3 T 166.122 73.5487 20170302_215400 -1 F F
20170302_221141 0.5,0.4,0.3 T 170.125 76.01 20170302_215400 -1 F F
20170302_221554 0.5,0.4,0.3 T 155.299 69.3328 20170302_221400 -1 F F
20170302_221606 0.5,0.4,0.3 T 154.173 69.2181 20170302_221400 -1 F F
20170302_221617 0.5,0.4,0.3 T 153.611 69.7213 20170302_221400 -1 F F

We increase the flux and we try the old recs with 1.0" disturbance.
TN AO Gain AO CL TT Res. WF [nm rms] HO Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20170302_222318 0.5,0.5,0.5 T 146.467 64.1705 20161013_153800 -1 F F
20170302_222331 0.5,0.5,0.5 T 136.967 63.3609 20161013_153800 -1 F F
20170302_222344 0.5,0.5,0.5 T 145.323 63.0442 20161013_153800 -1 F F
20170302_222714 0.5,0.5,0.4 T 128.111 61.164 20161013_153800 -1 F F
20170302_222951 0.5,0.5,0.5 T 143.1 67.4517 20161013_183000 -1 F F
20170302_223003 0.5,0.5,0.5 T 148.177 67.905 20161013_183000 -1 F F
20170302_223015 0.5,0.5,0.5 T 152.616 71.4606 20161013_183000 -1 F F
20170302_223342 0.5,0.5,0.4 T 138.747 65.2475 20161013_183000 -1 F F

When I try to run the old modal history in Open Loop the ASM rip (stroke too big?).

-- TommasoMazzoni - 03 Mar 2017
Topic revision: r1 - 03 Mar 2017, TommasoMazzoni
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback